# Scheme-independent determination of the QCD running coupling at all scales from jet observables using the principle of maximum conformality and infinite-order scale setting

### v3

@@ IV.1 Comment on the BFR method and results
The BFR strategy is based on the assumption that all bins are independent and statistically valid. Upon this hypothesis there is no risk of overfitting or of overestimating the errors or of biassing for both the central values and the errors. The BFR method leads to a bin interval where the theoretical perturbative predictions are more realistic and thus fit better the data. In this region nonperturbative effects and hadronization corrections are low in agreement with the range selected in Ref. [4]. Moreover, the BFR strategy tends to minimizing the discrepancy between theoretical predictions and data, by determining the minimum $\chi^{2}_{\mathrm{min}}$, and to maximizing the $n_{\rm dof}$ degrees of freedom, by the lowest $P_{\chi^{2}}(\chi^{2}_{\mathrm{min}};n_{\rm dof})$ and no other extra criteria is assumed to set the BFR interval. The lowest $P_{\chi^{2}}(\chi^{2}_{\mathrm{min}};n_{\rm dof})$ leads to the interval, that with the highest probability, we can consider the BFR leading to the best fit for the strong coupling in agreement with the Maximum-Likelihood strategy.
@@ V Conclusion
We have also introduced a method to improve the statistics consistent with the Maximum-Likelihood approach. The basic result of this method is to determine the best fitting region (BFR), i.e. the interval of bins that leads to the highest statistics with respect to all possible intervals of bins. By applying this method, the PMC and using the solution at 5-loop accuracy for the strong coupling, to thrust and $C$-parameter, we have obtained results that are consistent with the perturbative predictions and the extracted values of the strong coupling at the $Z^{0}$ mass scale, are consistent with the world average and with a significantly improved precision with respect to previous results.
@@ Acknowledgements.
LDG thanks Matthias Steinhauser for helpful support on the RunDec code. XGW is supported in part by the Natural Science Foundation of China under Grant No.12175025 and No.12347101. SQW is supported in part by the Natural Science Foundation of China under Grant No. 12265011. SJB is supported in part by the Department of Energy Contract No. DE-AC02-76SF00515. SLAC-PUB-17782.
Experiment $(1{-}T)$ – BFR $\alpha_{s}(\mathbf{M_{Z}})$ Tot. Err. $\chi^{2}_{\mathbf{min}}/n_{\mathrm{dof}}$ $P_{\chi^{2}_{\mathbf{min}}}^{n_{\mathrm{dof}}}$ $C$ – BFR $\alpha_{s}(\mathbf{M_{Z}})$ Tot. Err. $\chi^{2}_{\mathbf{min}}/n_{\mathrm{dof}}$ $P_{\chi^{2}_{\mathbf{min}}}^{n_{\mathrm{dof}}}$
ALEPH $(0.10;0.22)$ $0.1170$ ${}^{+0.0012}_{-0.0010}$ 0.80/11 $1.6\cdot 10^{-5}$ $(0.40;0.74)$ $0.1181$ ${}^{+0.0013}_{-0.0009}$ 2.63/16 $6.8\cdot 10^{-5}$
OPAL $(0.10;0.22)$ $0.1186$ ${}^{+0.0015}_{-0.0011}$ 5.0/2 0.92 $(0.40;0.74)$ $0.1187$ ${}^{+0.0016}_{-0.0011}$ 2.79/2 $0.75$
DELPHI $(0.10;0.22)$ $0.1191$ ${}^{+0.0012}_{-0.0008}$ 21.9/5 $0.9995$ $(0.40;0.74)$ $0.1198$ ${}^{+0.0015}_{-0.0010}$ 94.6/8 $\sim 1.0$
L3 $(0.10;0.22)$ $0.1175$ ${}^{+0.0011}_{-0.0009}$ 2.17/4 0.30 $(0.40;0.74)$ $0.1180$ ${}^{+0.0013}_{-0.0008}$ 6.86/6 0.67
Table 1: Results for the value of the strong coupling $\alpha_{s}(M_{Z})$ at 5-loop accuracy as implemented in the RunDec program, using the PMC for the thrust and the $C$-parameter distributions at NNLO and using OPAL, DELPHI and L3 data [24, 25, 23] in the BFR fixed using ALEPH [22].
@@
   $0.220.10<(1{-}T)<$ 0.22 (BFR) $0.1170$ ${}^{+0.0012}_{-0.0010}$ ${}^{+0.0010}_{-0.0006}$ ${}^{+0.0007}_{-0.0007}$ 0.80/11 $1.6\cdot 10^{-5}$
   $0.420.02<(1{-}T)<$ 0.42 (TFR) $0.1167$ ${}^{+0.0016}_{-0.0011}$ ${}^{+0.0016}_{-0.0010}$ ${}^{+0.0003}_{-0.0004}$ $195.1/39$ $\sim 1.0$
   $0.250.09<(1{-}T)<$ 0.25 (GFR) $0.1170$ ${}^{+0.0012}_{-0.0009}$ ${}^{+0.0010}_{-0.0006}$ ${}^{+0.0007}_{-0.0007}$ $3.9/15$ $1.9\cdot 10^{-3}$
Table 2: Results according to different interval ranges for the value of the strong coupling $\alpha_{s}(M_{Z})$ at 5-loop accuracy, as implemented in the RunDec program, using the PMC for the Thrust, $(1{-}T)$-distribution, at NNLO.
@@
   $0.220.04<(1{-}T)<$ 0.22 (BFR) $0.1170$ ${}^{+0.0018}_{-0.0011}$ ${}^{+0.0017}_{-0.0010}$ ${}^{+0.0005}_{-0.0004}$ 2.0/17 $2.9\cdot 10^{-6}$
   $0.420.02<(1{-}T)<$ 0.42 (TFR) $0.1177$ ${}^{+0.0017}_{-0.0012}$ ${}^{+0.0017}_{-0.0011}$ ${}^{+0.0003}_{-0.0003}$ $186.5/39$ $\sim 1.0$
   $0.250.09<(1{-}T)<$ 0.25 (GFR) $0.1174$ ${}^{+0.0013}_{-0.0009}$ ${}^{+0.0011}_{-0.0006}$ ${}^{+0.0007}_{-0.0006}$ $3.0/15$ $4.2\cdot 10^{-4}$
Table 3: Results according to different intervals for the value of the strong coupling $\alpha_{s}(M_{Z})$ at 5-loop accuracy, using the perturbative formula Eq. (I) and the PMC for the Thrust, $(1{-}T)$-distribution, at NNLO.
@@
   $0.260.09<(1{-}T)<$ 0.26 (BFR) $0.1285$ ${}^{+0.0059}_{-0.0042}$ ${}^{+0.0058}_{-0.0041}$ ${}^{+0.0008}_{-0.0009}$ 2.2/16 $1.8\cdot 10^{-5}$
   $0.420.02<(1{-}T)<$ 0.42 (TFR) $0.1327$ ${}^{+0.0059}_{-0.0039}$ ${}^{+0.0059}_{-0.0039}$ ${}^{+0.0006}_{-0.0005}$ $240.8/39$ $\sim 1.0$
   $0.250.09<(1{-}T)<$ 0.25 (GFR) $0.1285$ ${}^{+0.0060}_{-0.0041}$ ${}^{+0.0059}_{-0.0040}$ ${}^{+0.0009}_{-0.0008}$ $2.0/15$ $2.9\cdot 10^{-5}$
Table 4: Results according to different interval ranges for the value $\alpha_{s}(M_{Z})$ of the strong coupling at 5-loop accuracy as implemented in the RunDec program, using the conventional scale setting for the Thrust, $(1{-}T)$-distribution, at NNLO.

### v4

@@ IV.1 Comment on the BFR method and results
The BFR is designed to eliminate the errors arising from the bin-interval selection in the perturbative region, which have been neglected in the previous analyses. This method leads to a bin interval in which the theoretical perturbative predictions better fit the data. In this region non-perturbative effects and hadronization corrections are lower, in agreement with the range selected in Ref. [4]. Moreover, the BFR strategy tends to minimize the discrepancy between theoretical predictions and data by determining the minimum $\chi^{2}_{\mathrm{min}}$, and maximizing the $n_{\rm dof}$ the number of degrees of freedom, by the lowest $P_{\chi^{2}}(\chi^{2}_{\mathrm{min}};n_{\rm dof})$; no extra criteria are assumed to set the BFR interval. The lowest $P_{\chi^{2}}(\chi^{2}_{\mathrm{min}};n_{\rm dof})$ leads to the interval, that with the highest probability, we can consider the BFR leading to the best fit for the strong coupling in agreement with the Maximum-Likelihood strategy.
This strategy is based on the assumption that all bins are independent and statistically valid and is recommended when the errors given by the assumption of a wrong theoretical model are larger than the statistical errors. When this condition is satisfied, there is, in principle, no risk of overfitting or of overestimating the errors or of biassing either the central values or errors. If the statistical fluctuations are comparable with the theoretical model discrepancies, there could be a risk of selection bias; this is particularly evident when the BFR-interval is, for instance, in the non-perturbative region and it could be subsequently discarded, but it would not be detectable when the BFR is in the perturbative region. However, we expect that the latter case would occur when a bin-interval with a low number of bins is selected (e.g. when $n_{\rm dof}<5$) leading to both poor statistics and a not-too-good $\chi^{2}_{\rm min}$ value. In general, given that the BFR tends to maximise the number of bins in the interval, we would expect that in the case of a good statistical sample, this would be quite rare to occur.
@@ V Conclusion
We have also proposed a new method to eliminate the uncertainties related to the choice of the interval in the perturbative region, which is consistent with the Maximum-Likelihood approach. The purpose of this method is to determine the best fitting region (BFR), i.e. the interval of bins in the perturbative region, that best fit the data among all possible intervals of bins. By applying the BFR method, the PMC and using the strong coupling at 5-loop accuracy for thrust and the $C$-parameter, we have obtained results that are consistent with the perturbative predictions and the extracted values of the strong coupling at the $Z^{0}$ mass scale are consistent with the world average but with improved precision with respect to previous results. For further investigations, we would recommend using the BFR strategy as a refining strategy, i.e. as a methodology to be applied once there is sufficient knowledge of the theoretical predictions and we also would recommend further investigation of the method itself by improving either the treatment of systematic errors and correlation effects or by performing a “look-elsewhere” analysis in order to test the methodology and its results to the highest degree.
@@ Acknowledgements.
LDG thanks Matthias Steinhauser for helpful support on the RunDec code. XGW is supported in part by the Natural Science Foundation of China under Grant No.12175025 and No.12347101. SQW is supported in part by the Natural Science Foundation of China under Grant No. 12265011; by the Project of Guizhou Provincial Department under Grant No.YQK[2023]016 and No. ZK[2023]141. SJB is supported in part by the Department of Energy Contract No. DE-AC02-76SF00515. SLAC-PUB-17782.
Experiment (1T)(1{-}T) – BFR $\alpha_{s}(\mathbf{M_{Z}})$ Tot. Err. Expt. Err. Th. Err. $\chi^{2}_{\mathbf{min}}/n_{\mathrm{dof}}$ $P_{\chi^{2}_{\mathbf{min}}}^{n_{\mathrm{dof}}}$
ALEPH $(0.10;0.22)$ $0.1170$ $^{+0.0025}_{-0.0024}$ $^{+0.0023}_{-0.0023}$ $^{+0.0010}_{-0.0006}$ 0.80/11 $1.6\cdot 10^{-5}$
OPAL $(0.10;0.22)$ $0.1186$ $^{+0.0018}_{-0.0015}$ $^{+0.0014}_{-0.0014}$ $^{+0.0012}_{-0.0007}$ 5.0/2 0.92
DELPHI $(0.10;0.22)$ $0.1191$ $^{+0.0013}_{-0.0010}$ $^{+0.0007}_{-0.0007}$ $^{+0.0011}_{-0.0007}$ 21.9/5 $0.9995$
L3 $(0.10;0.22)$ $0.1175$ $^{+0.0013}_{-0.0010}$ $^{+0.0009}_{-0.0009}$ $^{+0.0010}_{-0.0006}$ 2.17/4 0.30
Table 1: Results for the value of the strong coupling $\alpha_{s}(M_{Z})$ at 5-loop accuracy as implemented in the RunDec program, using the PMC for thrust distribution at NNLO and using OPAL, DELPHI and L3 data [24, 25, 23] in the BFR fixed using ALEPH [22].
Experiment $C$ – BFR $\alpha_{s}(\mathbf{M_{Z}})$ Tot. Err. Expt. Err. Th. Err. $\chi^{2}_{\mathbf{min}}/n_{\mathrm{dof}}$ $P_{\chi^{2}_{\mathbf{min}}}^{n_{\mathrm{dof}}}$
ALEPH $(0.40;0.74)$ $0.1181$ $^{+0.0023}_{-0.0022}$ $^{+0.0020}_{-0.0020}$ $^{+0.0011}_{-0.0007}$ 2.63/16 $6.8\cdot 10^{-5}$
OPAL $(0.40;0.74)$ $0.1187$ $^{+0.0016}_{-0.0013}$ $^{+0.0011}_{-0.0011}$ $^{+0.0012}_{-0.0006}$ 2.79/2 $0.75$
DELPHI $(0.40;0.74)$ $0.1198$ $^{+0.0018}_{-0.0014}$ $^{+0.0010}_{-0.0010}$ $^{+0.0015}_{-0.0009}$ 94.6/8 $\sim 1.0$
L3 $(0.40;0.74)$ $0.1180$ $^{+0.0013}_{-0.0008}$ $^{+0.0006}_{-0.0006}$ $^{+0.0011}_{-0.0006}$ 6.86/6 0.67
Table 2: Results for the value of the strong coupling $\alpha_{s}(M_{Z})$ at 5-loop accuracy as implemented in the RunDec program, using the PMC for $C$-parameter distribution at NNLO and using OPAL, DELPHI and L3 data [24, 25, 23] in the BFR fixed using ALEPH [22].
@@
   0.10<(1T)<0.220.10<(1{-}T)<0.22 (BFR) $0.1170$ $^{+0.0012}_{-0.0010}$ $^{+0.0010}_{-0.0006}$ $^{+0.0007}_{-0.0007}$ 0.80/11 $1.6\cdot 10^{-5}$
   0.02<(1T)<0.420.02<(1{-}T)<0.42 (TFR) $0.1167$ $^{+0.0016}_{-0.0011}$ $^{+0.0016}_{-0.0010}$ $^{+0.0003}_{-0.0004}$ $195.1/39$ $\sim 1.0$
   0.09<(1T)<0.250.09<(1{-}T)<0.25 (GFR) $0.1170$ $^{+0.0012}_{-0.0009}$ $^{+0.0010}_{-0.0006}$ $^{+0.0007}_{-0.0007}$ $3.9/15$ $1.9\cdot 10^{-3}$
Table 3: Results according to different interval ranges for the value of the strong coupling $\alpha_{s}(M_{Z})$ at 5-loop accuracy, as implemented in the RunDec program, using the PMC for the Thrust, (1T)(1{-}T)-distribution, at NNLO.
@@
   0.04<(1T)<0.220.04<(1{-}T)<0.22 (BFR) $0.1170$ $^{+0.0018}_{-0.0011}$ $^{+0.0017}_{-0.0010}$ $^{+0.0005}_{-0.0004}$ 2.0/17 $2.9\cdot 10^{-6}$
   0.02<(1T)<0.420.02<(1{-}T)<0.42 (TFR) $0.1177$ $^{+0.0017}_{-0.0012}$ $^{+0.0017}_{-0.0011}$ $^{+0.0003}_{-0.0003}$ $186.5/39$ $\sim 1.0$
   0.09<(1T)<0.250.09<(1{-}T)<0.25 (GFR) $0.1174$ $^{+0.0013}_{-0.0009}$ $^{+0.0011}_{-0.0006}$ $^{+0.0007}_{-0.0006}$ $3.0/15$ $4.2\cdot 10^{-4}$
Table 4: Results according to different intervals for the value of the strong coupling $\alpha_{s}(M_{Z})$ at 5-loop accuracy, using the perturbative formula Eq. (I) and the PMC for the Thrust, (1T)(1{-}T)-distribution, at NNLO.
@@
   0.09<(1T)<0.260.09<(1{-}T)<0.26 (BFR) $0.1285$ $^{+0.0059}_{-0.0042}$ $^{+0.0058}_{-0.0041}$ $^{+0.0008}_{-0.0009}$ 2.2/16 $1.8\cdot 10^{-5}$
   0.02<(1T)<0.420.02<(1{-}T)<0.42 (TFR) $0.1327$ $^{+0.0059}_{-0.0039}$ $^{+0.0059}_{-0.0039}$ $^{+0.0006}_{-0.0005}$ $240.8/39$ $\sim 1.0$
   0.09<(1T)<0.250.09<(1{-}T)<0.25 (GFR) $0.1285$ $^{+0.0060}_{-0.0041}$ $^{+0.0059}_{-0.0040}$ $^{+0.0009}_{-0.0008}$ $2.0/15$ $2.9\cdot 10^{-5}$
Table 5: Results according to different interval ranges for the value $\alpha_{s}(M_{Z})$ of the strong coupling at 5-loop accuracy as implemented in the RunDec program, using the conventional scale setting for the Thrust, (1T)(1{-}T)-distribution, at NNLO.